# QUITE: A Query Rewrite System Beyond Rules with LLM Agents

### v2

@@ 7.3. Ablation Study
Table 8. Robustness Study of Different Data Sizes
Methods TPC-H (SF=1) TPC-H (SF=10) TPC-H (SF=30)
Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
Original 29.48 0.69 1.20 300.00 69.31 9.42 30.08 300.00 64.76 20.61 46.69 300.00
LearnedRewrite 1.46 0.52 1.10 3.91 37.57 9.97 30.55 202.31 52.59 18.90 43.50 300.00
LLM-R2(Deepseek-R1) 15.38 0.69 1.22 7.69 56.73 9.25 29.33 300.00 54.67 18.30 58.11 300.00
R-Bot (GPT-4o) 1.49 0.66 1.03 6.38 32.64 9.48 26.25 172.20 54.54 19.85 61.12 300.00
QUITE 1.02 0.64 1.01 3.79 26.06 9.28 24.19 75.64 40.65 16.89 37.96 131.97
Table 9. Data-Aware Analysis on TPC-H Benchmark
Methods TPC-H (SF=1) TPC-H (SF=10) TPC-H (SF=30)
Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
Original 29.48 0.69 1.20 300.00 69.31 9.42 30.08 300.00 64.76 20.61 46.69 300.00
QUITE (Rewrite on SF=1) 1.02 0.64 1.01 3.79 36.87 13.53 29.02 279.51 54.36 25.02 43.03 289.48
QUITE (Rewrite on SF=10) 1.08 0.68 1.22 3.79 26.06 9.28 24.19 75.64 45.90 17.44 57.26 155.72
QUITE (Rewrite on SF=30) 5.72 0.50 1.17 5.17 30.50 9.96 24.29 96.68 40.65 16.89 37.96 131.97
@@
Query Hint Injection. We evaluate the impact of hint injection on queries rewritten by QUITE. As shown in Table 3, the insertion of hints leads to a further decrease in average execution times by 1.8%, 3.8%, and 0.3% on TPC-H, DSB, and Calcite, respectively. These performance gains demonstrate that appropriate hints enable additional fine-grained optimization, even when queries have already been extensively optimized through rewriting. To highlight the advantage of our LLM-based hint injection approach, we compare its performance on rewritten queries against using LLMs to recommend Bao’s candidate hint sets. As shown in Table 7 and Figure 7, within the hint space defined at the GUC level provided by Bao, the overall trend indicates degraded performance.
Furthermore, Figure 7 presents a comparison of execution times for 24 positively optimized queries on DSB, both before and after hint injection, as well as against the original queries. Notably, some queries that exhibited degraded performance after rewriting were able to exceed their original execution efficiency through proper hint injection. This demonstrates that hint injection can effectively recover and enhance performance in cases where the initial rewrites were suboptimal.
Base Model Selection.To evaluate the role of the base models within our framework, we conducted experiments using different model combinations. DeepSeek-R1 was retained as the Reasoning agent due to its low cost and strong chain-of-thought reasoning capabilities, making it indispensable for this role. For the Decision agent, we tested three alternatives: Claude-3.7-Sonnet, DeepSeek-V3, and GPT-4o. As shown in Table 6, rewrites using Claude-3.7-Sonnet as the Decision agent significantly outperform those from the other two models. This is due to its strengths in summarization and decision-making, which require long-context understanding, precise reasoning, and accuracy—areas where Claude-3.7-Sonnet excels. In contrast, DeepSeek-V3 shows higher hallucination rates, and GPT-4o offers lower accuracy and weaker long-context handling, making both less suitable for our framework.
@@
LLM-R2 (Claude-3.7) 399.30 0.93
R-Bot (Claude-3.7) 256.96 20.94
QUITE (DS-R1 + Claude-3.7) 291.59 6.76
@@ 7.4. Robustness Study
We assess the robustness of QUITE by testing stability across data scales. We scale the TPC-H dataset from SF = 10 to SF = 1 and SF = 30, then execute rewritten queries on all three versions. Table 8 shows that QUITE delivers significant optimizations at each scale and outperforms all baselines. In contrast, LearnedRewrite and LLM-R2 often produce rewrites that worsen performance for complex queries.
We evaluate the robustness of the QUITE framework by analyzing its stability. The stability assessment demonstrates that QUITE generalizes effectively across diverse data scales. To evaluate the stability of QUITE across varying database sizes, we extended the primary TPC-H $SF=10$ dataset by generating proportionally scaled versions at scale factors of $SF=1$ and $SF=30$. For each evaluated method, queries rewritten based on the $SF=10$ dataset were executed and timed on the $SF=1$, $SF=10$, and $SF=30$ datasets. As shown in Table 8, QUITE consistently achieved substantial optimization improvements across all scales, outperforming all baseline methods.
@@ 7.5. Further Analysis
Data-Aware Analysis. We evaluated the sensitivity of QUITE using a cross-testing strategy. Specifically, QUITE generated three sets of rewritten queries, each optimized for the data statistics of a specific TPC-H scale: $SF$=1, $SF$=10, and $SF$=30. Each set was subsequently executed across all three datasets. As shown in Table 9, the best performance consistently occurred when queries were executed on the dataset for which they were originally optimized. This result confirms that QUITE effectively distinguishes between different data distributions and produces rewrites that are well-adapted to the target datasets.
Cost Analysis. As shown in Table 10, LLM-R2 incurs high time expenditure due to demonstration construction, while R-Bot consumes substantial tokens during stepwise rewriting. In contrast, QUITE offers a balanced trade-off between time expenditure and cost. By generating much of its output with DeepSeek-R1, it greatly reduces monetary cost. Additionally, it avoids pre-training and extensive knowledge base construction. To further control execution time, we cap FSM iterations, ensuring reliable output quality without incurring excessive time expenditure.
@@ 8. Conclusion
This paper presents QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance. We rethink the query rewrite process and identify unique opportunities to support a wider range of query patterns and rewrite strategies with good scalability. The implemented QUITE  system significantly outperforms state-of-the-art methods in both query performance and query coverage on widely used benchmarks.

### v3

@@ 7.3. Ablation Study
Table 7. Robustness Study on TPC-H and DSB Benchmarks under Different Scale Factors (SF)
Methods TPC-H DSB
SF=1 SF=10 SF=30 SF=1 SF=10 SF=30
Mean Med 75th 95th Mean Med 75th 95th Mean Med 75th 95th Mean Med 75th 95th Mean Med 75th 95th Mean Med 75th 95th
Original 29.48 0.69 1.20 300.00 69.31 9.42 30.08 300.00 64.76 20.61 46.69 300.00 7.22 0.27 0.69 13.19 32.62 4.85 10.29 300.00 39.42 5.06 18.22 300.00
LearnedRewrite 1.46 0.52 1.10 3.91 37.57 9.97 30.55 202.31 52.59 18.90 43.50 300.00 17.15 0.42 1.52 81.14 31.93 5.14 15.18 251.42 43.88 7.09 20.73 300.00
LLM-R2(DS-R1) 15.38 0.69 1.22 7.69 56.73 9.25 29.33 300.00 54.67 18.30 58.11 300.00 4.30 0.23 0.57 2.15 9.16 3.84 8.21 24.21 23.01 5.01 11.87 87.15
R-Bot (GPT-4o) 1.49 0.66 1.03 6.38 32.64 9.48 26.25 172.20 54.54 19.85 61.12 300.00 9.06 0.25 0.63 21.89 21.75 4.26 8.80 119.19 31.93 5.15 16.27 300.00
QUITE (Rw. on SF=1) 1.02 0.64 1.01 3.79 36.87 13.53 29.02 279.51 54.36 25.02 43.03 289.48 2.25 0.24 0.50 2.74 10.58 4.18 8.59 30.79 23.23 4.67 12.91 111.83
QUITE (Rw. on SF=10) 1.08 0.68 1.22 3.79 26.06 9.28 24.19 75.64 45.90 17.44 57.26 155.72 2.82 0.24 0.51 1.72 6.08 3.78 7.98 21.43 17.36 4.34 9.28 77.62
QUITE (Rw. on SF=30) 5.72 0.50 1.17 5.17 30.50 9.96 24.29 96.68 40.65 16.89 37.96 131.97 6.65 0.26 0.56 3.20 8.74 3.94 8.01 23.15 16.35 4.45 9.26 74.33
LLM-Agent-based Query Rewrite FSM. To evaluate the impact of the FSM framework, we make two modifications: (1) remove the equivalence-checking loop and (2) remove the decision stage. This converts the process into a linear pipeline while keeping all components and agent tools for rewriting. To further isolate the contribution of the MDP-based reasoning agent, we also conduct an ablation that removes the MDP component. As shown in Table 6, QUITE outperforms all variants across metrics. Notably, adding the MDP reduces the average execution time by 52.2% compared to the version without it, indicating that the MDP substantially improves the reasoning agent’s effectiveness.
@@
Query Hint Injection. We evaluate the impact of hint injection on queries rewritten by QUITE. As shown in Table 4, hint injection further reduces average execution time by 1.8%, 3.8%, and 0.3% on TPC-H, DSB, and Calcite, respectively, indicating that appropriate hints enable additional fine-grained optimization even after extensive rewriting. To highlight the advantage of our approach, we compare it with using LLMs to select Bao’s GUC-level hint sets. As shown in Table 5 and Figure 8, this baseline exhibits an overall performance degradation within Bao’s predefined hint space. Furthermore, Figure 8 shows execution times for 24 positively optimized DSB queries before and after hint injection, as well as for the original queries. In several cases, hint injection recovers and surpasses the original performance when rewriting alone is suboptimal. This demonstrates hint injection can recover and enhance performance in cases where the initial rewrites were suboptimal.
Figure 10. Rewrite Categories for Different Benchmarks
@@
LLM-R2 (Claude-3.7) 391.23 0.016
R-Bot (Claude-3.7) 118.33 0.417
QUITE (DS-R1 + Claude-3.7) 218.25 0.200
Sensitivity Study. We examine QUITE’s sensitivity to the number of rewrite rounds, hint thresholds, and different LLM models. We first test different model combinations. We fix DeepSeek-R1 as the reasoning model and vary the other models among Claude-3.7-Sonnet, DeepSeek-V3, and GPT-4o, and then we replace the reasoning model with DeepSeek-V3 and Claude-3.7-Sonnet. As shown in Table 5, the combination of DeepSeek-R1’s strong reasoning capabilities and Claude-3.7-Sonnet’s long-context understanding significantly outperforms all other combinations.
For hint thresholds, as shown in Table 5, we test the effect of adding the scan method to our hint base and the effect of using all hints (excluding GUC). The former yields an execution time nearly identical to that of the rewritten query, while the latter causes a slight performance degradation. This suggests that the size of our current hint search space is appropriate.
@@ 7.4. Robustness Study
Scale-Aware Robustness. We evaluate QUITE on TPC-H and DSB under scale factors $SF\in\{1,10,30\}$, where rewriting and execution are performed on the same scale. As Table 7 shows, QUITE consistently achieves substantial latency reductions at every scale. In contrast, LR, LLM-R2, and R‑Bot often degrade as the data scale changes. These results demonstrate that QUITE maintains stable efficiency under varying data volumes and adapts robustly to distribution shifts within each benchmark.
Cross-Scale Robustness. To examine generalization across data distributions, we perform cross-scale evaluation: rewrites produced at one scale ( $SF\in\{1,10,30\}$) are executed on all others. As Table 7 shows, QUITE achieves its strongest gains when rewrite and execution scales match, while mismatched cases show only minor regressions and still outperform all baselines. This behavior highlights QUITE’s strong data adaptivity and robustness, effectively learning to tailor rewrites to underlying data characteristics without overfitting to a specific workload or scale.
@@ 7.5. Further Analysis
Rewrite Type Analysis. We report successful rewrites that achieve over a 10% performance improvement while maintaining equivalence across four benchmarks. As shown in Figure 10, most gains come from join, CTE, and predicate rewrites, while the rest are benchmark-specific. Compared with rule-based baselines (lr; llm-r2; r-bot), QUITE produces deeper rewrites beyond pattern matching. We further analyse efficient rewrite types, query structures, Q&A units, and query hints in our Github files (quite_ap; quite_rt).
Cost Analysis. We conducted a cost analysis experiment on 63 queries of TPC-H. As shown in Table 8, LLM-R2 incurs high time cost for demonstration construction, while R-Bot consumes substantial tokens for stepwise rewriting. In contrast, QUITE offers a balanced trade-off between time expenditure and cost. By generating much of its output with DeepSeek-R1, QUITE greatly reduces monetary cost. Additionally, it avoids pre-training and extensive knowledge base construction.
@@ 8. Conclusion
This paper presents QUITE, a training-free and feedback-aware system that smartly leverages LLM Agents to rewrite SQL queries into semantically equivalent forms with significantly improved performance. We rethink the query rewrite process and identify unique opportunities to support a wider range of query patterns and rewrite strategies with strong scalability. The implemented QUITE  system significantly outperforms state-of-the-art methods in both query performance and query coverage on widely used benchmarks and synthetic workloads.